1. We are given the initial pressure of 1.5 atm and temperature of 25⁰C of the gas and asked for the pressure at a higher temperature of 450⁰C 
1
Expert's answer
2021-10-27T06:41:14-0400

Let's assume that the volume of gas remains unchanged.

P1 = 1.5 atm

T1 = 25⁰C = 298 K

V1 = V2 = constant

P2 = unknown

T2 = 450⁰C = 723 K


Solution:

Since the volume doesn't change, Gay-Lussac's law can be used.

Gay-Lussac's law can be expressed as: P1/T1 = P2/T2

P1T2 = P2T1

To find the final pressure, solve the equation for P2:

P2 = P1T2/T1

P2 = (1.5 atm × 723 K) / (298 K) = 3.6393 atm = 3.6 atm

P2 = 3.6 atm


Answer: The final pressure of gas is 3.6 atm